Solve the equation 2x3 -9x2+12x-4 =0 given that two of its roots are equal
Answer by lwsshak3(11628)

Solve the equation 2x3 -9x2+12x-4 =0 given that two of its roots are equal.
***
Use rational roots theorem to solve:
...0...|.....2.....-9.......12......-4
...1...|.....2.....-7.......5.........1 (root betweeen 0 and 1
...2..|......2.....-5.......2.........0 (2 is a root)
f(x)=(x-2)(2x^2-5x+2)
f(x)=(x-2)(x-2)(2x-1)
roots are: 2, 2, 1/2